# Better Fit: Accommodate Variations in Clothing Types for Virtual Try-on

### v1

@@ Abstract
Keywords:
Virtual Try-On, Mask Inpainting, Try-On Evaluation
@@
1 Introduction
Image-based virtual try-on is a hot topic in the field of conditional image generation, aiming to create an image of a person wearing a target clothing item given the clothing image and the person image. It offers practical advantages for online shoppers and opens up new avenues for creativity in the fashion domain.
@@ 1 Introduction
The key reason of changing clothing types attributes to that the try-on area is affected by the original clothing that is desired to take off. Forcing constraints in clothing warping [33] could keep target clothing away from excessive distortion, but also hinders flexible deformation particularly when facing challenging postures. Paying attention to the width-height ratio of clothing and perform truncation [35] could avoid squeezing clothing, but cannot deal well with various clothing types. We propose a novel adaptive mask training paradigm to break the correlation between the try-on area and the original clothing, which corrects the bias caused by paired training data, thus training the model to learn more accurate semantic correspondences for inpainting.
As shown in the right-most column of Fig. 1, the proposed strategy significantly enhances the performance of unpaired try-on.
@@
2 Related Work
2.1 Image Based Virtual Try-On
Early virtual try-on methods[12, 29, 21, 33, 4, 19, 8, 13] are based on GANs[31]. These methods divide virtual try-on into two parts: clothing warping[20, 7, 17] and the fusion of warped clothing with the model. After Diffusion[6] model demonstrated stronger image generation capabilities than GANs, works[22, 10] use pretrained Diffusion models[26, 32] as the fusion module to generate final try-on images. However, limited by the performance of the clothing warping module, when model pose are complex, errors in clothing warping can propagate to the final try-on results. Recent methods[36, 18] seek to accomplish virtual try-on solely using Mask Inpainting Diffusion models, directly generating the clothing and body in the try-on state within the mask area while preserving the content outside the mask. Existing methods do not consider the differences in try-on between training and inference, leading to the model erroneously interpreting the lower boundary of the mask area as the lower boundary of the target clothing, as illustrated in Fig. 2. We successfully addressed this flaw by applying adaptive masks for the training samples during the training process.
2.2 Metrics For Virtual Ty-On
@@
3 Adaptive Mask Inpainting for Virtual Try-On
@@
3.1 Adaptive Mask Maker
@@
For checking boundary, We first utilize the openpose[3] parser to obtain three key points around the waist, and then interpolate an additional key point on each side, resulting in a total of five points used as checkpoints. The preliminary judgment on whether the bottom interfere with the top is made by checking if these five points fall within the semantic area[9] of the top. If more than $\tau_{B}$ checkpoints are in the area of the top, the wearing style is classified as Interfered. Otherwise, we need to proceed to the next step for further analysis.
@@
Wearing Style={ $\text{Wearing Style}=\begin{cases}\text{\leavevmode\nobreak\ Non-Interfered},&\text{Checkpoints}>\tau_{B}\\ \text{\leavevmode\nobreak\ Non-Interfered},&\text{Checkpoints}\leq\tau_{B},\leavevmode\nobreak\ R_{\text{T}}\geq\tau_{T}\\ \text{\leavevmode\nobreak\ Interfered},&\text{Checkpoints}\leq\tau_{B},\leavevmode\nobreak\ R_{\text{T}}<\tau_{T}\end{cases}$ (1)
@@
3.2 Adaptive Mask Training Paradigm
@@
4 Evaluation Metrics for Unpaired Try-On
@@
4.1 Semantic Densepose Ratio
@@
We introduce correction factors utilizing the overlapping area between $S$ and $D$, denoted as $S\cap D$. $\alpha=\frac{D}{S\cap D}$ represents the fabric area factor, which can amplify the distance between the SDR values of two results when the clothing has a type that uses less fabric, thereby capturing more subtle differences. $\beta=\frac{S\cap D}{S}$ represents the clothing fit factor, which can reduce the distance of SDR values when the clothing has a looser fit, thus providing a more lenient evaluation for looser type. In the testing of unpair try-on, we calculate the $\alpha$ and $\beta$ factors using $S_{R}$ and $D_{R}$ from the real try-on. The similarity metric between the Virtual try-on result and the Real try-on result, based on the SDR, can be represented by the equation 4. Our experiments indicate that the SDR distance can effectively measure the accuracy of clothing type in unpaired try-on, with more results referenced in Section 5.
$SDR\ Distance=\frac{D_{R}}{S_{R}\cap D_{R}}\cdot\frac{S_{R}\cap D_{R}}{S_{R}}\cdot\left|\frac{S_{R}}{D_{R}}-\frac{S_{V}}{D_{V}}\right|=\left|1-\frac{D_{R}S_{V}}{S_{R}D_{V}}\right|$ (4)
4.2 Skeleton Based LPIPS
@@
5 Experiment
5.1 Experimental Setup
Datasets. We tested our training paradigm and evaluation metrics on VITON-HD[4]. The VITON-HD dataset contains 14,221 training samples and 2,032 testing samples. We conduct unpaired try-on experiments on two test sets: Unpair-2032 and Cross-27. Unpair-2032 is the original test set of VITON-HD, composed of random try-on combinations of 2032 test samples. Cross-27 is a manually curated cross-try-on benchmark. We selected 27 samples from the VITON-HD test set by considering three dimensions: long sleeves / short sleeves / sleeveless, long/short lengths, and Interfered / Non-Interfered. We performed cross-try-on among these samples to obtain 729 try-on groups. Our training and testing were both conducted at a resolution of $512\times 384$.
@@ 5.1 Experimental Setup
5.2 Mask Effectiveness Analysis
@@
VITON-HD HR-VTON LaDI-VTON DCI-VTON GP-VTON StableVITON Ours
Figure 8: Comparison results on Unpair-2032.
Figure 9: Comparison results on Cross-27.
Table 1: Quantitative comparison results on Unpair-2032 and Cross-27.
@@
Quantitative analysis. We conducted quantitative analysis using the Unpair-2032 and our Cross-27. Since MGD[1] has not open-sourced its sketch parser, we are unable to obtain the test results of MGD on the Cross-27. As of the writing of this manuscript, StableVITON[18] has not provided training code, so we replicated the code based on the information provided by the paper, it is denoted as Baseline in Table 1. On the Unpair-2032, our training paradigm inherits the advantages of StableVITON and achieves better SDR and S-LPIPS scores compared to Baseline. Although our FID and KID scores are worse than those of some previous works, this does not imply that our try-on quality is similarly inferior; we will explain the reasons for this occurrence in the section 5.3. The effectiveness of our method is further demonstrated on Cross-27, where, compared to Baseline, we lead significantly in SDR and S-LPIPS scores.
Ablation analysis. To verify the effectiveness of our adaptive mask training paradigm, we conducted training with three different masks in addition to our adaptive mask training paradigm. Small Mask: following the mask approach in DCI-VITON[10], we inflate the semantic area of the top clothing to serve as the mask; Big Mask: following the method in TryOnDiffusion[36], we use a rectangular area below the face and above the waist as the mask; Normal Mask: we adopt the same mask approach as StableVITON[18]. As shown in Fig. 10, only our training paradigm effectively addresses the flaw of the lower boundary of the top clothing, preserving the type of the target clothing. Table 2 presents the quantitative comparison. The adaptive mask paradigm maintains the same advantages in the SDR and S-LPIPS metrics as seen in the visualization results. We will explain in the section 5.3 why the FID and KID scores for the adp-mask under Cross-27 testing are inferior to those of other masks.
Figure 10: Comparison results from training with different masks.
Table 2: Quantitative comparison results on different masks.
@@
Table 3: Comparison of scores with different proportions of Incorrect samples.
@@
5.3 Analysis of Metric Effectiveness
@@
We conducted experiments on both Unpair-2032 and Cross-27, mixing Incorrect samples into the unpaired results at intervals of 20%, with the calculation results of all metrics shown in Table 3. To more intuitively reflect the sensitivity of each metric to Incorrect samples, we normalized the score and plotted it in Fig. 11. FID and KID both fail to accurately reflect the negative impact of Incorrect samples on try-on results. The inclusion of Incorrect samples actually makes the data distribution of try-on results closer to that of the test dataset, leading to a trend in FID that is contrary to the actual situation on both test sets. Compared to FID, KID’s stronger and more precise comparison capability does not benefit the evaluation of try-on results but rather results in chaotic outcomes on the Unpair-2032. Overall, try-on results that tend to preserve the original clothing information of the model paradoxically achieve lower FID and KID scores (indicating better performance) which is clearly incorrect. Our proposed SDR and S-LPIPS metrics consistently reflected the impact of incorrect samples on try-on quality under all experimental conditions, significantly outperforming FID and KID in evaluating the accuracy of unpaired try-on.
6 Conclusions

### v2

@@ Abstract
Index Terms:
Image Generation, Virtual Try-On, Mask Inpainting, Try-On Evaluation
I Introduction
Recently, fashion images attract attentions [1, 2] and image-based virtual try-on is a hot topic in the field of conditional image generation, aiming to create an image of a person wearing a target clothing item given the clothing image and the person image. It offers practical advantages for online shoppers and opens up new avenues for creativity in the fashion domain.
@@ I Introduction
The key reason of changing clothing types attributes to that the try-on area is affected by the original clothing that is desired to take off. Forcing constraints in clothing warping [6] could keep target clothing away from excessive distortion, but also hinders flexible deformation particularly when facing challenging postures. Paying attention to the width-height ratio of clothing and perform truncation [10] could avoid squeezing clothing, but cannot deal well with various clothing types. We propose a novel adaptive mask training paradigm to break the correlation between the try-on area and the original clothing, which corrects the bias caused by paired training data, thus training the model to learn more accurate semantic correspondences for inpainting. As shown in the right-most column of Fig. 2, the proposed strategy significantly enhances the performance of unpaired try-on.
@@
II Related Work
II-A Image Based Virtual Try-On
Human-centered conditional image generation can be classified according to conditions, such as the text modality [25] and the image modality (e.g., pose [26]). Image-based virtual try-on takes the target clothing image as a condition and also the model image as input, which aims to synthesize an image where the model naturally wears target clothes. Early virtual try-on methods[3, 4, 5, 6, 8, 9, 7, 27] are based on GANs[28]. These methods divide virtual try-on into two parts: clothing warping[29, 30, 31] and the fusion of warped clothing with the model. After Diffusion[32] model demonstrated stronger image generation capabilities than GANs, works[16, 17] use pretrained Diffusion models[15, 33] as the fusion module to generate final try-on images. However, limited by the performance of the clothing warping module, when model pose are complex, errors in clothing warping can propagate to the final try-on results. Recent methods[34, 11] seek to accomplish virtual try-on solely using Mask Inpainting Diffusion models, directly generating the clothing and body in the try-on state within the mask area while preserving the content outside the mask. Existing methods do not consider the differences in try-on between training and inference, leading to the model erroneously interpreting the lower boundary of the mask area as the lower boundary of the target clothing, as illustrated in Fig. 1. We successfully addressed this flaw by applying adaptive masks for the training samples during the training process.
II-B Metrics For Virtual Ty-On
@@
III Adaptive Mask Inpainting for Virtual Try-On
@@
III-A Adaptive Mask Maker
@@
For checking boundary, We first utilize the openpose[37] parser to obtain three key points around the waist, and then interpolate an additional key point on each side, resulting in a total of five points used as checkpoints. The preliminary judgment on whether the bottom interfere with the top is made by checking if these five points fall within the semantic area[38] of the top. If more than $\tau_{B}$ checkpoints ($cp$) are in the area of the top, the wearing style is classified as Interfered. Otherwise, we need to proceed to the next step for further analysis.
@@
Wearing Style={ $\text{Wearing Style}=\begin{cases}\text{\leavevmode\nobreak\ Non-Interfered},&\text{cp}>\tau_{B}\\ \text{\leavevmode\nobreak\ Non-Interfered},&\text{cp}\leq\tau_{B},\leavevmode\nobreak\ R_{\text{T}}\geq\tau_{T}\\ \text{\leavevmode\nobreak\ Interfered},&\text{cp}\leq\tau_{B},\leavevmode\nobreak\ R_{\text{T}}<\tau_{T}\end{cases}$ (1)
@@
III-B Adaptive Mask Training Paradigm
@@
IV Evaluation Metrics for Unpaired Try-On
@@
IV-A Semantic Densepose Ratio
@@
We introduce correction factors utilizing the overlapping area between $S$ and $D$, denoted as $S\cap D$. $\alpha=\frac{D}{S\cap D}$ represents the fabric area factor, which can amplify the distance between the SDR values of two results when the clothing has a type that uses less fabric, thereby capturing more subtle differences. $\beta=\frac{S\cap D}{S}$ represents the clothing fit factor, which can reduce the distance of SDR values when the clothing has a looser fit, thus providing a more lenient evaluation for looser type. In the testing of unpair try-on, we calculate the $\alpha$ and $\beta$ factors using $S_{R}$ and $D_{R}$ from the real try-on. The similarity metric between the Virtual try-on result and the Real try-on result, based on the SDR, can be represented by the equation 4. Our experiments indicate that the SDR distance can effectively measure the accuracy of clothing type in unpaired try-on, with more results referenced in Section V.
$\begin{split}SDR\ Distance&=\frac{D_{R}}{S_{R}\cap D_{R}}\cdot\frac{S_{R}\cap D_{R}}{S_{R}}\cdot\left|\frac{S_{R}}{D_{R}}-\frac{S_{V}}{D_{V}}\right|\\ &=\left|1-\frac{D_{R}S_{V}}{S_{R}D_{V}}\right|\end{split}$ (4)
IV-B Skeleton Based LPIPS
@@
V Experiment
V-A Experimental Setup
Datasets. We tested our training paradigm and evaluation metrics on VITON-HD [8]. The VITON-HD dataset contains 14,221 training samples and 2,032 testing samples. We conduct unpaired try-on experiments on two test sets: Unpair-2032 and Cross-27. Unpair-2032 is the original test set of VITON-HD, composed of random try-on combinations of 2032 test samples. Cross-27 is a manually curated cross-try-on benchmark. We selected 27 samples from the VITON-HD test set by considering three dimensions: long sleeves / short sleeves / sleeveless, long/short lengths, and Interfered / Non-Interfered. Specifically, for each condition, we select three samples by considering different body shapes and poses. The selected samples in Cross-27 are shown in Fig. 8, and each person image corresponds to a clothing image. Cross-27 Benchmark is used for evaluating the cross-try-on performance with various and challenging try-on conditions, and 729 (i.e., $27\times 27$) try-on images in total are obtained. Our training and testing were both conducted at a resolution of $512\times 384$.
Non-Interfered Interfered
Long Short
Sleeveless Short Sleeves Long Sleeves Sleeveless Short Sleeves Long Sleeves Sleeveless Short Sleeves Long Sleeves
Figure 8: All samples in Cross-27.
Fig. 9 compares the distribution of try-on situations covered by Unpair-2032 and Cross-27, where Cross-27 exhibits a more balanced distribution. Try-on situations in blue and orange are less challenging, resulting in a bias on easy samples for VITON-HD.
Figure 9: Distribution of try-on situations in test set.
@@ V-A Experimental Setup
V-B Mask Effectiveness Analysis
@@
Table I: Quantitative comparison results on Unpair-2032 and Cross-27.
@@
Table II: Quantitative comparison results on different masks.
@@
VITON-HD HR-VTON LaDI-VTON DCI-VTON GP-VTON StableVITON Ours
Figure 10: Comparison results on Unpair-2032.
Quantitative analysis. We conducted quantitative analysis using the Unpair-2032 and our Cross-27. Since MGD[18] has not open-sourced its sketch parser, we are unable to obtain the test results of MGD on the Cross-27. As of the writing of this manuscript, StableVITON[11] has not provided training code, so we replicated the code based on the information provided by the paper, it is denoted as Baseline in Table I. On the Unpair-2032, our training paradigm inherits the advantages of StableVITON and achieves better SDR and S-LPIPS scores compared to Baseline. Although our FID and KID scores are worse than those of some previous works, this does not imply that our try-on quality is similarly inferior; we will explain the reasons for this occurrence in the section V-C. The effectiveness of our method is further demonstrated on Cross-27, where, compared to Baseline, we lead significantly in SDR and S-LPIPS scores.
Ablation analysis. To verify the effectiveness of our adaptive mask training paradigm, we conducted training with three different masks in addition to our adaptive mask training paradigm. Small Mask: following the mask approach in DCI-VITON[17], we inflate the semantic area of the top clothing to serve as the mask; Big Mask: following the method in TryOnDiffusion[34], we use a rectangular area below the face and above the waist as the mask; Normal Mask: we adopt the same mask approach as StableVITON[11]. As shown in Fig. 12, only our training paradigm effectively addresses the flaw of the lower boundary of the top clothing, preserving the type of the target clothing. Table II presents the quantitative comparison. The adaptive mask paradigm maintains the same advantages in the SDR and S-LPIPS metrics as seen in the visualization results. We will explain in the section V-C why the FID and KID scores for the adp-mask under Cross-27 testing are inferior to those of other masks.
Figure 11: Comparison results on Cross-27.
Figure 12: Comparison results from training with different masks.
Figure 13: Visual analysis of SDR and S-LPIPS metrics.
(a) Unpair-2032
(b) Cross-27
Table III: Comparison of scores with different proportions of Incorrect samples.
@@
V-C Analysis of Metric Effectiveness
@@
We conducted experiments on both Unpair-2032 and Cross-27, mixing Incorrect samples into the unpaired results at intervals of 20%, with the calculation results of all metrics shown in Table III. To more intuitively reflect the sensitivity of each metric to Incorrect samples, we normalized the score and plotted it in Fig. 15. FID and KID both fail to accurately reflect the negative impact of Incorrect samples on try-on results. The inclusion of Incorrect samples actually makes the data distribution of try-on results closer to that of the test dataset, leading to a trend in FID that is contrary to the actual situation on both test sets. Compared to FID, KID’s stronger and more precise comparison capability does not benefit the evaluation of try-on results but rather results in chaotic outcomes on the Unpair-2032. Overall, try-on results that tend to preserve the original clothing information of the model paradoxically achieve lower FID and KID scores (indicating better performance) which is clearly incorrect. Our proposed SDR and S-LPIPS metrics consistently reflected the impact of incorrect samples on try-on quality under all experimental conditions, significantly outperforming FID and KID in evaluating the accuracy of unpaired try-on.
(a) FID
(b) KID
(c) S-LPIPS
(d) SDR
Figure 15: Comparison of FID, KID, S-LPIPS, SDR scores, and negative user preference rates for each method.
Figure 16: Comparison results of the same model trying on different clothing.
Figure 17: Comparison results of different models trying on the same clothing.
Fig. 13 shows two examples containing both good fit and not good fit results, where the try-on quality is also reflected by the proposed metrics. The main advantage in try-on results against previous methods is the preservation of clothing types, which is measured by SDR. The try-on situation that demands highly in type preservation is marked with green in Fig. 9, which occupies a small portion in VITON-HD. Therefore, our results acquire marginal improvement on Unpair-2032, but significant improvement in SDR on Cross-27. Besides comparing clothing types by SDR, S-LPIPS is proposed to compare clothing texture, both of which are proposed to make up the drawbacks of existing metrics in unpaired evaluation.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 18: Our method achieves realistic visual effects when trying on various types of clothing.
V-D User Study
We conducted a survey with 40 human users using the results generated on the Cross-27. Among the try-on results of the seven methods (in Fig. 15), we randomly selected the try-on results of three methods each time. To reflect the relative performance of seven methods, users were asked to choose the "most realistic" results among three randomly selected samples instead of seven. If making one choice from seven methods, we probably get one or two recent methods with higher scores while the others with poor performance. Each user completed 50 such choices. We calculate the user preference rate (%) of each method as the proportion of times that it was chosen as "most realistic" out of the total number of choices. A higher percentage indicates that users consider the results generated by that method to be more realistic.
Since the scores of the metrics are negatively correlated with the quality of the try-on, for a more convenient comparison, we take the negative of the user preference rate to obtain the "Negative User Preference Rate." As shown by the red line in Fig. 15, our method achieved the lowest negative user preference rate, which means the highest user preference rate, among all methods. In addition, when comparing the scores of FID, KID, S-LPIPS, and SDR metrics with the negative user preference rate, the relationship of superiority and inferiority among methods reflected by FID and KID shows a considerable deviation from the user preference. In contrast, our proposed S-LPIPS and SDR metrics are more consistent with the user preference, as shown in Fig. 15(c) and Fig. 15(d). SDR is designed to reflect the extent of changing clothing type via try-on, with amplification for poor results. However, user evaluation assigns the same weight for each testing case. Therefore, compared with user study, the performance of seven methods on SDR experiences fluctuation, but shares a similar trend (e.g., the high score of user study becomes much higher for SDR at DCI-VTON and StableVITON while the low score of user study becomes much lower for SDR at GP-VTON and Ours.
V-E More Visual Results
We present more comparison results in Fig. 16 and Fig. 17, where our method shows significant improvements in preserving the type of clothing compared to existing work. Our method is well compatible with trying on clothing of different types, as shown in Fig. 18.
V-F Limitations & Discussion
Although our method can well maintain the type of the clothing, when the model’s original clothing is relatively cumbersome (covering a larger area), our mask cannot completely eliminate the original clothing. These residual clothing pixels will appear as flaws in the final generated results, as shown within the red box in Fig. 19. In our subsequent work, we will further explore more refined methods of mask creation to eliminate all original clothing information while preserving the non-try-on parts.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 19: The model’s original clothing is too cumbersome to be completely masked, resulting in undesired remaining in the generated images.
The proposed evaluation metrics also rely on parsing. We give examples of parsing flaws in Fig. 20, where Densepose contains holes due to the color of background/foreground, cloth segment is too long due to the same color of top/bottom garment, and pose detector misses some keypoints. As failed parsing cases occupy a very small portion, the training is not affected with abundant training samples. With the help of corrections in our implementation, e.g., 1) Densepose and semantic segmentation cooperate with each other to fill holes and 2) keypoints fitering as illustrated in Sec. IV-B, the limitations on failed parsing can be alleviated, but not totally overcome.
Figure 20: Limitation: parsing flaws
VI Conclusions